OpenStudy (anonymous):

One person can do a certain job in ten minutes, and another person can do the same job in fifteen minutes. How many minutes will they take to do the job together?

OpenStudy (anonymous):

7.5 minutes

OpenStudy (anonymous):

that isnt one of the answers the options are A) 4 min b) 5 min c) 6 min and please show me the process

OpenStudy (anonymous):

i apologise it is 4 minutes, i mis calculated. so if we call 10 minute man A and 15 minute man B. Lets take the job as having 15 little tasks. A does 1.5 tasks/minute B does 1 task/ minute. make sense ?

OpenStudy (anonymous):

then together they do 2.5 tasks per minute and 15/2.5 is not 4 but 6 haha wrong twice

OpenStudy (anonymous):

The formula for this one isn't too hard. Add the rates, keeping in mind that the rate is the reciprocal of the time. For this problem,\[\frac{1}{10}+\frac{1}{15}=\frac{1}{T}\]where T is the time working together.

OpenStudy (anonymous):

i think the total time is 25 if everybody do the work himself then 10/25 +15/25

OpenStudy (anonymous):

\[\frac{1}{10}+\frac{1}{15}=\frac{1}{T} \implies \frac{5}{30}=\frac{1}{T}\implies T=6\]

OpenStudy (anonymous):

ok that makes sense thanks

OpenStudy (anonymous):

Keep working. Do math every day.